# Physics-Aware, Shannon-Optimal Compression via Arithmetic Coding for Distributional Fidelity

### v1

@@ III.3 Lossless Compression as a Likelihood-Scored Fidelity Diagnostic
Arithmetic coding provides an operational realization of negative log-likelihood under a fixed reference model. Let $q_{A}$ denote the probability model (CDF tables) learned from a reference dataset $A$ and then held fixed. For a dataset $D=\{x_{i}\}_{i=1}^{N}$ encoded with this codec, the total codelength satisfies
@@
Here $\hat{p}_{D}$ denotes the empirical distribution associated with dataset $D$, while the unknown population distribution is denoted by $p$. Given an independent baseline sample $B$ from the same population as $A$ and a perturbed (or synthetic) sample $C_{\varepsilon}$, we define the excess codelength
@@ IV.1 Invertibility
To verify the lossless and invertible nature of the proposed arithmetic coding framework, we perform a direct comparison between detector-level observables computed from the original data and from the data obtained after a full compression–decompression cycle. The dataset $R$ is split into disjoint training ($A$) and validation ($B$) subsets, with models trained on $A$ and evaluated on $B$. Figure 1 reports representative results for dataset $B$.
@@ IV.3 Comparison between Shannon Entropy and Average Sequence Length
Tables 2 and 3 report the empirical entropy $H(q)$, the cross-entropy $H(p,q)$, and the achieved average code length $L$ (all in bits per event) for the unconditional and conditional codecs, respectively. Results are shown separately for the calorimeter hits and for the particle-kinematics payload, as well as for the total event representation, and for both the training dataset $A$ and the validation $B$. Several features are worth highlighting.
@@
Second, the cross-entropy evaluated on the validation dataset $B$ is systematically larger than that obtained on the training dataset $A$. This increase is consistent with the expected additional “surprise” incurred when encoding previously unseen data drawn from the same underlying process, reflecting finite-sample estimation effects rather than a true distributional difference, since the reference model is trained on $A$ rather than on $B$ [18].
@@ IV.5 Fidelity Studies
Let $q_{A}$ denote the probabilistic model (CDFs) learned from the reference sample $A$ and then held fixed. Following Eqs. (11) and (12), for a dataset $D=\{x_{i}\}_{i=1}^{N}$ the arithmetic codec induces an average per-event codelength (in bits per event), $L_{A}(D)$, up to negligible termination overhead (Sec. IV.3). We define the excess codelength
@@
evaluated under the same fixed reference codec trained on $A$. Let $\mu_{C}(\varepsilon)$ and $\mu_{B}$ denote the sample means of $\{L_{A}(C_{\varepsilon,k})\}$ and $\{L_{A}(B_{k})\}$, respectively, with sample standard deviations $s_{C}(\varepsilon)$ and $s_{B}$. We define the mean excess codelength

### v2

@@ III.3 Lossless Compression as a Likelihood-Scored Fidelity Diagnostic
Arithmetic coding provides an operational realization of negative log-likelihood under a fixed reference model. Let $q_{A}$ denote the probability model (CDF tables) learned from a reference dataset AA and then held fixed. For a dataset $D=\{x_{i}\}_{i=1}^{N}$ encoded with this codec, the total codelength satisfies
@@
Here $\hat{p}_{D}$ denotes the empirical distribution associated with dataset $D$, while the unknown population distribution is denoted by pp. Given an independent baseline sample $B$ from the same population as AA and a perturbed (or synthetic) sample $C_{\varepsilon}$, we define the excess codelength
@@ IV.1 Invertibility
To verify the lossless and invertible nature of the proposed arithmetic coding framework, we perform a direct comparison between detector-level observables computed from the original data and from the data obtained after a full compression–decompression cycle. The dataset $R$ is split into disjoint training (AA) and validation ($B$) subsets, with models trained on AA and evaluated on $B$. Figure 1 reports representative results for dataset $B$.
@@ IV.3 Comparison between Shannon Entropy and Average Sequence Length
Tables 2 and 3 report the empirical entropy $H(q)$, the cross-entropy $H(p,q)$, and the achieved average code length $L$ (all in bits per event) for the unconditional and conditional codecs, respectively. Results are shown separately for the calorimeter hits and for the particle-kinematics payload, as well as for the total event representation, and for both the training dataset AA and the validation $B$. Several features are worth highlighting.
@@
Second, the cross-entropy evaluated on the validation dataset $B$ is systematically larger than that obtained on the training dataset AA. This increase is consistent with the expected additional “surprise” incurred when encoding previously unseen data drawn from the same underlying process, reflecting finite-sample estimation effects rather than a true distributional difference, since the reference model is trained on AA rather than on $B$ [18].
@@ IV.5 Fidelity Studies
Let $q_{A}$ denote the probabilistic model (CDFs) learned from the reference sample AA and then held fixed. Following Eqs. (11) and (12), for a dataset $D=\{x_{i}\}_{i=1}^{N}$ the arithmetic codec induces an average per-event codelength (in bits per event), $L_{A}(D)$, up to negligible termination overhead (Sec. IV.3). We define the excess codelength
@@
evaluated under the same fixed reference codec trained on AA. Let $\mu_{C}(\varepsilon)$ and $\mu_{B}$ denote the sample means of $\{L_{A}(C_{\varepsilon,k})\}$ and $\{L_{A}(B_{k})\}$, respectively, with sample standard deviations $s_{C}(\varepsilon)$ and $s_{B}$. We define the mean excess codelength